OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? 1300 1120 708 412

OpenStudy (anonymous):

SOMEONE PLEASE HELP.

OpenStudy (radar):

Read the question and convert to equivalent equations. First thing label the two type of equations. As an example, label the one-step equation as 'x' and the two-step equation as 'y', There is no hard fast rule for the labels you could of labeled the one step as OS and the two-step as TS ( but I am digressing)

OpenStudy (radar):

we have been informed by the problem that a total of the equations (including both types) is 1120. Converting this to algebra talk we write: x + y = 1120 Equation 1. We have also been informed that there is a algebraic relationship between the amount of x and y. Converting this to math speak we get: 3x - 2y = 1300 Equation 2. Before going any further, is there any question regarding the converting the written word to algebra?

OpenStudy (radar):

This is not going to be a monolog, earth to sammiegurl3, come in please.

OpenStudy (anonymous):

hi sorry.

OpenStudy (radar):

Have you reviewed my two post and do you understand what we have accomplished?

OpenStudy (radar):

You now have two equations with only two unknowns (x and y). This is a solvable situation using substitution or elimination. I am leaving it up to you to choose the method you prefer. You asked for help, but you are not helping yourself by your silence.

OpenStudy (radar):

so I ask you do you really need help?

OpenStudy (radar):

x + y =1120 3x-2y=1300 Solve for x (one-step equations) Solve for y (two-step equations)

OpenStudy (anonymous):

i dont know how to start if i have only variables.

OpenStudy (radar):

What methods have you been studying in order to solve linear equations with two variables?

OpenStudy (radar):

There are two most common methods. One is "substitution" and the other is "elimination". Have you ever heard of these methods?

OpenStudy (radar):

I suggest you review your instructional material. Here is how I would solve them. I would add the two equations getting: 4x-y=2420 I would then add the first equation to this: x +y=1120 ------------- 5x = 3540 Solve for x by dividing by 5 X=708 That is the number of one-step BUT it asked for the number of two step equation using x+y=1120 708+y=1120 y=1120-708=412 That is your answer. 412 two step equation. Once again learn how to solve, then if you need help, then it will be helpful.

OpenStudy (radar):

Open study is not for giving you answers, but to work with you in the solution.

OpenStudy (radar):

Good luck with these kind of problems.

OpenStudy (anonymous):

thanks.